### v1

# Know Your Needs Better: Towards Structured Understanding of Marketer Demands with Analogical Reasoning Augmented LLMs

Abstract.
In this paper, we explore a new way for user targeting, where non-expert marketers could select their target users solely given demands in natural language form. The key to this issue is how to transform natural languages into practical structured logical languages, i.e., the structured understanding of marketer demands. Considering the impressive natural language processing ability of large language models (LLMs), we try to leverage LLMs to solve this issue. Past research indicates that the reasoning ability of LLMs can be effectively enhanced through chain-of-thought (CoT) prompting. But existing methods still have some limitations: (1) Previous methods either use simple “Let’s think step by step” spells or provide fixed examples in demonstrations without considering compatibility between prompts and questions, making LLMs ineffective in some complex reasoning tasks such as structured language transformation. (2) Previous methods are often implemented in closed-source models or excessively large models, which is not suitable in industrial practical scenarios. Based on these, we propose ARALLM (i.e., Analogical Reasoning Augmented Large Language Models) consisting of two modules: Analogical Reasoning based Prompting and Reasoning-Augmented Multi-Task Model Distillation. To be specific, we first introduce a structured language SELL that is practical in marketing scenarios and construct a reasoning library consisting of several high-quality reasoning examples of translating natural language to SELL (NL2SELL). Then, we adopt a retrieval-based method to conduct analogical reasoning with the help of the reasoning library. The experimental results show that prompting with analogical reasoning achieves better performance than the ordinary prompting method in NL2SELL tasks. Beyond that, we distill knowledge from super LLMs (ChatGPT) based on analogical reasoning and construct a NL2SELL dataset to fine-tune smaller student LLMs in a multi-task training paradigm, enabling the models to be easily deployed in practical user targeting environments.
User Targeting, Large Language Models, Analogical Reasoning, Knowledge Distillation, Chain of Thought
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
Figure 1. The comparison of user targeting approach.
1. Introduction
Recently, the practice of user targeting has gained significant attention in real-world applications (e.g., Alipay and WeChat), as highlighted by a variety of studies (Yang et al., 2023a; Mangalampalli et al., 2011; Shen et al., 2015; Zhuang et al., 2020; Zhu et al., 2021). This approach has the excellent potential to attract high-quality users for specific campaigns, aligning with the goals of marketers to enhance conversions and minimize operational expenses. Roughly speaking, current methods devoted to user targeting mainly fall into two lines: model-based methods (Figure 1 (a)) that perform expansion of seed users/entities with well-designed neural networks (Yang et al., 2023a; Zhuang et al., 2020; Zhu et al., 2021) and rule-based methods (Figure 1 (b)) that manually group users with different tags based on domain knowledge (Mangalampalli et al., 2011; Shen et al., 2015).
Unfortunately, both of the above prior studies have primarily emphasized the intricate architectures for effectively and efficiently gathering targeted users, while commonly ignoring the natural and significant gap between marketers’ demand and the capability of current models. In particular, current user targeting pipelines mainly force the marketers to decompose their demands into multiple tags/entities, e.g., tags such as “Preference”, “Resident City” and “Age” are characterized in Figure 1. The process incurs several weaknesses: i) unfriendly to marketers; ii) time-consuming for manual deconstruction of demands; and iii) unpromising performance due to partial tags/entities.
Therefore, in this paper, we take the initial stride towards enhancing current user targeting systems by probing into the structured understanding of marketers’ demands in an automatic way, enabling the system to know your needs better solely based on natural language inputs. As shown in Figure 1 (c), the marketers only need to input their native demand like “Young people in Hangzhou who enjoy milk tea or white-collar workers in first-tier cities who enjoy listening to audiobooks” and our new approach can instantly provide the user targeting cards through a novelty structured understanding of the demand. After verifying the cards, they can export the results with just one click and the system will automatically run the existing targeted user aggregation modules. The whole process only takes 3-4 minutes, which is 10 times faster than the former approach. The core of this new approach lies in how to transform natural languages into structured forms, or so-called structured understanding of marketer demands.
To address this issue, we first develop a Structured and Editable Logical Language called SELL, which forms the basis of the user targeting cards shown in Figure 1(c). The main concern of our work is transforming the Natural Language into SELL (NL2SELL). For example, the natural language “Young people in Hangzhou who enjoy milk tea or white-collar workers in first-tier cities who enjoy listening to audiobook” can be translated into SELL as “((Resident City#Belongs To# Hangzhou) AND (Age#Between#18,35) AND (Preference#Belongs To#Milk Tea)) OR ((City Level#Belongs To#First-tier) AND (Days of listening to audiobooks#Greater Than#3) AND (Career#Belongs To#White-collar))”. Previous research  (Liu et al., 2023; Gu et al., 2023; Yang et al., 2023b; Lu et al., 2022) has highlighted the remarkable abilities of Large Language Models (LLMs) in language translation tasks. LLMs are renowned for their robust natural language processing and zero-shot capabilities, which render them possibly effective in NL2SELL. Consequently, we opt to employ LLMs to tackle this task.
However, there are still some challenges in using LLMs in our scenario: (1) The challenge of reasoning accuracy. There is still room for improvement in reasoning accuracy for LLMs in complex problems, such as structured understanding of natural language and translation issues in our scenario. Recently, it (Wei et al., 2022; Kojima et al., 2022) has been pointed out that prompting LLMs with chain-of-thought (CoT) could enhance the reasoning ability of LLMs. These methods either provide fixed reasoning examples as demonstrations through few-shot learning or tell the model “Let’s think step by step” for zero-shot learning. Though effective, they ignore the compatibility between prompts and specific questions. In some challenging reasoning scenarios, zero-shot learning or having few related reasoning examples in few-shot learning may not effectively stimulate the LLMs’ reasoning ability (Rubin et al., 2021). Ideally, we can customize a reasoning process for each question to induce the model to reason better, but it is time-consuming and laborious. (2) The challenge of reasoning speed and resource consumption. Although the reasoning ability of super-large language models can be developed through prompt engineering, it is unacceptable to deploy excessively huge LLMs or API-based LLMs in practical situations. For extremely huge LLMs, excessive parameters will consume a lot of resources, and the reasoning speed is also slower. For API-based LLMs such as ChatGPT, both the security of private data and the cost of API calls are thorny issues. Thus, smaller, free, and white-box LLMs are preferred.
To address these challenges, we propose a framework called Analogic Reasoning Augmented Large Language Model (ARALLM) to tackle the NL2SELL task, which consists of two modules: Analogical Reasoning based Prompting and Reasoning-Augmented Multi-Task Model Distillation. (1) Analogical Reasoning based Prompting. Creating a CoT for each new targeting demand is impractical and time-consuming (considering the need for deployment). Thus, we propose Analogical Reasoning based Prompting. The fundamental concept of it is that when two targeting demands are similar, they might follow similar reasoning steps and have similar conditional expression structures. Therefore, we construct a compact yet reliable reasoning library to serve as a reference for numerous demands that have unknown reasoning steps. This approach helps reduce the effort of manually writing reasoning steps. Additionally, we can offer assistance for any unfamiliar demands by employing analogical reasoning with retrieval. (2) Reasoning-Augmented Multi-Task Model Distillation. Since the white-box and smaller model is preferred, we first distill knowledge from teacher LLMs(ChatGPT) to construct a NL2SELL dataset with over 10,000 pieces of training samples and then propose a novel multi-task training method with reasoning augmentation.
We summarize our contributions as follows: (1) Aiming to automatically understand marketers’ demands in a structured manner, we propose a brand-new language SELL and NL2SELL task. (2) We propose an innovative framework called ARALLM, which is composed of the Analogical Reasoning based Prompting module and the Reasoning-Augmented Multi-Task Distillation module, to address the NL2SELL task. (3) We demonstrate the superiority of the proposed ARALLM framework through extensive experiments on real-world datasets.
2. Design of SELL
In this section, we will provide a detailed explanation of the design and the usage of the SELL. The SELL is mainly composed of four elements:
Keys are series of tags that describe the features of the user, such as Age, Gender, Monthly Income, Pet Owning, and so on. In practical marketing scenarios, the database contains tens of thousands of tags that describe user profiles. Selecting the appropriate keys from this vast array of tags is a challenging task.
Values are the fillings of the corresponding keys. They can be generally divided into three types: numerical type, string type, and boolean type. For instance, the value corresponding to the key Monthly Income is of numerical type, because they are continuous and non-enumerable. The value of Gender is of string type, with limited options: Male or Female. Boolean type values have only two states, i.e., True or False, such as Pet Owning.
Operators represent the relationship between keys and values. For numerical values, seven types of operators can be used to connect them with corresponding keys: Equal To, Greater Than, Less Than, Not Equal To, Not Greater Than, Not Less Than, and Between. For values of string type and boolean type, there are two types of operators: Belongs To and Not Belongs To. Keys, operators, and values can form a basic conditional expression in SELL, formatted as “(key#operator#value)”, such as “(Pet Owning#Belongs To#True)”. A conditional expression represents a cluster of target users, for example, in the above example, the target users are pet owners.
Intersection and Union symbols are used to take the intersection or union of multiple conditional expressions. In practical marketing scenarios, the features of the target users are often complex and difficult to describe using a single conditional expression. Therefore, it is necessary to combine the target users through intersection or union operations. In SELL, we use the symbols AND and OR to denote the intersection and union operations, respectively. For example, “((Resident City#Belongs To#Hangzhou) OR (Resident City#Belongs To#Shanghai)) AND (Pet Owning#Belongs To#True)” describes the pet owners who live in Hangzhou or Shanghai.
Compared to the popular database languages SQL, the advantages of SELL are: (1) SELL is simpler in expression and has much fewer syntax symbols than SQL, making it easier for non-programming experts to understand. (2) Since SELL can be transformed into a clear tree structure, marketers can easily modify them even if the generated results have flaws.
3. Method
Figure 2. Overview of analogical reasoning based prompting methods.
In this work, we focus on how to convert a sentence of marketer’s demand $d$ into a standard SELL expression, while how to parse and execute the SELL and return the target users from the database will be handled by the internal system, which does not constitute the core of this article. Inspired by the analogical reasoning (Yao et al., 2023; Turney, 2008; Gentner, 1983) in the classic artificial intelligence field and CoT prompting methods in LLMs, we first propose a novel prompting method enhanced by analogical reasoning and then adopt a reasoning-augmented multi-task training strategy to fine-tune smaller LLMs.
3.1. Prompting With Analogical Reasoning
The core idea of analogical reasoning in the NL2SELL task is that if two demands are similar, they may share similar reasoning steps, as well as similar conditional expression structures. Thus, we construct a small but high-quality reasoning library to provide references for numerous demands with unknown reasoning steps, minimizing the cost of manually writing reasoning steps for each of them.
3.1.1. Construction of Reasoning Library
To construct a reasoning library, we first randomly select 50 marketer demands from real-world scenarios and write their corresponding answers based on our expert knowledge to form Q&A pairs =$\mathcal{R}=\{(d_{e},s_{e})\}_{e=1}^{50}$, where $d_{e}$ is the demand, $s_{e}$ is the corresponding SELL expressions. Thanks to the simplicity of the SELL itself, this processing can be completed by two experts within an hour. Then we use a small amount of manual effort to write corresponding reasoning details for 5 different ($d_{e}$,$s_{e}$) $\in \mathcal{R}$ to construct the seed examples in the format $(d_{e},r_{e},s_{e})$, where $r_{e}$ represent the reasoning details. We generally summarize the procedure of solving the NL2SELL problem as four steps:
  1. (1)
    Extract keywords: Extract the keywords from the demand.
  2. (2)
    Select tags: Select the most relevant tags from the tag list based on the keyword as the key in SELL.
  3. (3)
    Form conditional expressions: Based on the selected keywords and contextual information, fill in the corresponding operators and values to form conditional expressions.
  4. (4)
    Combine: Combine conditional expressions using intersection or union symbols.
What we need to do is to fill in the details of these four steps.
For the remaining data in $\mathcal{R}$, we utilize the OpenAI gpt-3.5-turbo API [^1] https://platform.openai.com/docs/api-reference/chat to help us fill the corresponding reasoning steps, and our seed examples serve as demonstrations in the prompt. After simple verification, we harvest a batch of solid reasoning details to form a reasoning library =$\mathcal{R}=\{(d_{e},r_{e},s_{e})\}_{e=1}^{50}$.
3.1.2. Retrieval-Based Analogical Prompt Construction
After the construction of the reasoning library, we can provide references for unknown demands through analogical reasoning. As shown in Figure 2, given a marketer demand $d_{p}$ that needs to predict the corresponding SELL expression, we first retrieve similar demands $d_{e}$ in $\mathcal{R}$ based on their text embedding similarity:
(1) $sim(d_{p},d_{e})=cos(\textbf{E}(d_{p}),\textbf{E}(d_{e})),d_{e}\in\mathcal{R}$
where E is the embedding model which we use BGE [^2] https://huggingface.co/BAAI/bge-large-zh(Xiao et al., 2023; Zhang et al., 2023) off the shelf as it is state-of-the-art in Chinese text retrieval field. The top $k$ demands in $\mathcal{R}$ that get the highest similarity score when compared with $d_{p}$ will be fetched as analogical examples in this stage, along with their corresponding $r$ and $s$.
In addition, to help the LLMs better organize answers, we should also inform the LLMs of the tags in the marketing database as explicit knowledge. To address the issue of an excessive number of tags, we still use BGE as a retriever to retrieve the most relevant n tags of every demand from the marketing database to construct a part of the prompt. The top n tags that get the highest similarity scores when compared with demand will form a small-scale tag list $T=\{t_{x}\}_{x=1}^{n}$ to provide a reference for selecting keys in SELL.
Therefore, the basic organizational form of the final prompt $x_{p}$ for predicting is :
(2) $x_{p}=concat(\mathcal{I},\{(d_{e},T_{e},r_{e},s_{e})\}_{e=1}^{k},(d_{p},T_{p}))$
where $\mathcal{I}$ is the instruction which simply describes the task, k is the number of analogical examples, $T_{e}$ and $T_{p}$ are the most relevant tag lists of the $d_{e}$ and $d_{p}$ respectively.
Compared to the ordinary CoT method, our approach exhibits several main differences:
First, the NL2SELL task is more challenging. Unlike solving mathematical problems or common sense reasoning tasks, the training data for LLMs don’t inherently include the syntax of SELL. This highlights the importance of appropriate reasoning steps.
Second, we place a greater emphasis on the compatibility between prompts and questions. We believe that using analogical reasoning based prompting method can make LLMs more adaptable to specific problems, rather than relying on fixed few-shot examples, which could potentially limit the model’s performance.
3.2. Reasoning-Augmented Multi-Task Model Distillation
To train a private small model for deployment, we first distill knowledge from teacher LLM (ChatGPT) to construct a NL2SELL dataset with over 10,000 pieces of training samples and then propose a multi-task training method with reasoning augmentation.
3.2.1. Knowledge Distillation
We collect nearly 10,000 demands from our daily marketing environment. Following the analogical reasoning method mentioned above, we construct input prompts for each demand in the form of Equation (2), where analogical examples come from the previously constructed reasoning library. Subsequently, based on these prompts that incorporate analogical knowledge, we call the gpt-3.5-turbo API to generate SELL answers, which will then be further verified and corrected through manual crowdsourcing. We denote this part of data comes from demand to answer paradigm.
It can be seen that the above distillation method relies on a large number of native demands and tedious manual calibration. At the same time, due to the different frequencies of user tags used in actual marketing activities, the distilled data has a long tail distribution. This means that some demands based on unpopular user tags only exist in a small amount of data, resulting in poor response to such demands. To address the above two issues, we propose a distillation approach based on answer to demand, which can be further divided into two stages: answer generation and demand generation.
$\bullet$ Answer generation: As shown in Table 5, we first construct a series of logical templates of SELL, where c is the basic conditional expression in SELL. When constructing an answer, we first randomly select an answer template from the logical template table and then construct the content of every c in the selected template to obtain a synthesized answer. The content of c is obtained from the marketing database through random sampling. Specifically, we first randomly sample a tag from the database as a key in c. Then we sample its operator based on the type of keys. Finally, we randomly select a value from all possible values of that key, thus constructing a c in the form of (key#operator#value). To ensure that the synthesized answers are logically reasonable, we will limit the number of conditional expressions in the answer template, and abandon overly complex intersection and union operations.
$\bullet$ Demand generation: Considering the strong writing ability of ChatGPT, we provide some ($d_{e},s_{e}$) examples in $\mathcal{R}$ as demonstrations to guide gpt-3.5-turbo to write demands given above synthesized answers:
(3) $concat((d_{e},s_{e}),s_{i})\xrightarrow{gpt-3.5-turbo}d_{i}$
where $s_{i}$ is the above synthesized SELL answers. We will check the language fluency and logical rationality of the generated demand, and the specific prompt content is shown in Appendix C.
We denote this part of data distilled through the answer to demand pipeline. We obtain 1,200 pieces of data in this way and combine them with data coming from the demand to answer pipeline to form the final training data $\mathcal{D}_{train}=\{(d_{i},s_{i})\}_{i=1}^{N}$ for small model, where N is the number of training data.
3.2.2. Reasoning-Augmented Multi-Task Fine-tune
The training objective of the normal distillation method is to minimize the loss L:
(4) $L=\frac{1}{N}\sum_{i=1}^{N}CE(f(d_{i}),s_{i}),(d_{i},s_{i})\in\mathcal{D}_{train}$
where f represents the model, $CE$ is the cross-entropy loss function.
It (Hsieh et al., 2023; Kim et al., 2023) has been pointed out that fine-tuning LLMs using data combined with reasoning steps can enhance the reasoning capabilities of the LLMs. Inspired by (Hsieh et al., 2023), we first prompt gpt-3.5-turbo to generate intermediate reasoning steps of $\mathcal{D}_{train}$ based on reasoning library $\mathcal{R}$:
(5) $concat(({d}_{e},r_{e},s_{e}),(d_{i},s_{i}))\xrightarrow{gpt-3.5-turbo}r_{i}$
where $(d_{e},r_{e},s_{e})\in\mathcal{R}$, $(d_{i},s_{i})\in\mathcal{D}_{train}$, $r_{i}$ is the corresponding generated reasoning steps of $(d_{i},s_{i})$. Subsequently, we define a multi-task training strategy in which not only $s_{i}$ but also $r_{i}$ are used as supervised labels to train the model.
To be specific, we denote the input of these two tasks as:
(6) $x_{i_{s}}=concat(\mathcal{I}_{s},\{(d_{e},T_{e},r_{e},s_{e})\}_{e=1}^{k},(d_{i},T_{i}))$
(7) $x_{i_{r}}=concat(\mathcal{I}_{r},\{(d_{e},T_{e},r_{e},s_{e})\}_{e=1}^{k},(d_{i},T_{i}))$
where the only difference between them is the content of instruction $\mathcal{I}_{s}$ and $\mathcal{I}_{r}$ which guide the model to generate the SELL answer or reasoning steps respectively. The corresponding training objectives of these two tasks are:
(8) $L_{s}=\frac{1}{N}\sum_{i=1}^{N}CE(f(x_{i_{s}}),s_{i})$
(9) $L_{r}=\frac{1}{N}\sum_{i=1}^{N}CE(f(x_{i_{r}}),r_{i})$
The final loss function of multi-task fine-tuning is then defined as:
(10) $L_{total}=L_{s}+L_{r}$
4. Experiments
Table 1. Data statistics. Number(d), MaxLen(d), MinLen(d) and AvgLen(d) denote the number, the maximum string length, the minimum string length and the average string length of the demands.
Number(d) MaxLen(d) MinLen(d) AvgLen(d)
$\mathcal{R}$ 50 42 7 16.5
$\mathcal{D}_{train}$ 11329 207 1 16.8
$\mathcal{D}_{test}$ 170 48 3 17.0
4.1. Experimental Settings
4.1.1. Benchmarks
We manually construct a NL2SELL testing benchmark for predicting, which contains 170 pieces of data based on expert knowledge in the format $\mathcal{D}_{test}=\{(d_{p},s_{p})\}_{p=1}^{170}$. The demands in $\mathcal{D}_{test}$ do not exist in the training set $\mathcal{D}_{train}$ or reasoning library $\mathcal{R}$, eliminating the issue of label leakage. Table 1 shows the statistics of these 3 parts of data [^3] The dataset does not contain any Personal Identifiable Information (PII). The dataset is desensitized and encrypted. Adequate data protection was carried out during the experiment to prevent the risk of data copy leakage..
4.1.2. Baseline
We first use ChatGPT to test the effectiveness of the analogical reasoning based prompting method. For comparison, we propose four additional variants of prompts as baselines:
$\bullet$ Zero-shot: Only given the instruction $\mathcal{I}$ and a test demand $d_{p}$.
$\bullet$ Fixed few-shot: Given the instruction $\mathcal{I}$, k fixed examples $\{(d_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for all $d_{p}\in\mathcal{D}_{test}$ and a $d_{p}$.
$\bullet$ Fixed few-shot + RS (Reasoning Steps): Given the instruction $\mathcal{I}$, k fixed examples $\{(d_{e},r_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for all $d_{p}\in\mathcal{D}_{test}$ and a $d_{p}$.
$\bullet$ Random few-shot + RS: Given the instruction $\mathcal{I}$, k randomly sampled examples $\{(d_{e},r_{e},s_{e})\}_{e=1}^{k}\in\mathcal{R}$ for every $d_{p}\in\mathcal{D}_{test}$ and a $d_{p}$. Different $d_{p}$ has different randomly sampled examples.
In the knowledge distillation stage, we choose ChatGLM2-6B-32K (Du et al., 2022) and Baichuan2-13B-Chat (Baichuan, 2023) as student LLMs, considering their better comprehension ability in dealing with long Chinese texts compared to other models. In addition to fine-tuning methods mentioned in Section 3.2.1, we also propose three other training strategies as baselines:
$\bullet$ - MT: Training LLMs without multi-task (MT) strategy. The model input is (6) and the training objective is (8).
$\bullet$ - RS: Training LLMs with multi-task strategy but the reasoning steps $r_{e}$ of examples in (6) and (7) are not given in the inputs, in which way we can explore the impact of analogical reasoning.
$\bullet$ Normal: Training LLMs in a most normal way, in which the input of the model is the simple combination of instruction $\mathcal{I}$ and a demand $d_{i}$, and the supervised label is $s_{i}$ only.
4.1.3. Evaluation metrics
To conduct a more reasonable and comprehensive evaluation, we evaluate the output of the LLMs on the NL2SELL task from 4 perspectives:
$\bullet$ Structure accuracy (Structure Acc.): The logic structure’s accuracy of the predicted SELL expression. We remove the keys, operators, and values in the SELL, leaving only intersection and union symbols and separators as logic structures. For example, the structure of “((Location#Belongs To#Hangzhou) OR (Location#Belongs To#Shanghai)) AND (Pet Owning#Belongs To#True)” is “((##) OR (##)) AND (##)”. We use the mean value of the Levenshtein distance (L) (Levenshtein et al., 1966) and the Ratcliff/Obershelp similarity (R/O) (Black, 2004) to evaluate the structure accuracy compared to label SELL.
$\bullet$ Overall accuracy (Overall Acc.): We use the ScareBLEU (Post, 2018) score to evaluate the overall similarity of the output and label texts as the overall accuracy. We set this as the main metric to evaluate the quality of predicted SELL answers.
$\bullet$ Human evaluation (Human Eval.): We conduct more fine-grained evaluations through human evaluation. Annotators who have received good training will score the accuracy of the key, value, and logic between 0 and 10, and finally provide an overall score. The accuracy of the key and value refers to whether the key and value in predicted answers are identical to or synonymous with labels. The accuracy of logic refers to whether the logic aligns with labels. Human evaluation is important because sometimes the model output may not be completely consistent with labels, but it may also be correct. For example, “(Age#Between#18,35)” could also be expressed as “(Age#Greater Than#18) AND (Age#Less Than#35)”, while the above evaluation metrics would treat them as different answers.
$\bullet$ GPT4 evaluation (GPT4 Eval.): We use the most advanced LLM “GPT4” as the judges. To be specific, we prompt the GPT4 to score the predicted answers from 0 to 10 and provide different scoring examples as a reference for GPT4 (as shown in Appendix D).
4.1.4. Implementation details
Considering the limitations of input length, we set the number of analogical examples k as 2. In the fine-tuning stage, ChatGLM2-6B-32K and Baichuan2-13B-Chat are tuned with LoRA (Hu et al., 2021) on 1 and 4 A100 (80G). We set the optimizer, maximum context size, batch size, and learning rate to Adam, 4096, 8, and 5e-5 respectively. For testing, a single 80G A100 is used. We use the code framework from LLaMA-Factory [^4] https://github.com/hiyouga/LLaMA-Factory (hiyouga, 2023).
Table 2. NL2SELL results on ChatGPT. The best results are in bold, while the second are underlined.
Overall Acc. Structure Acc. Human Eval. GPT4 Eval.
S-BLEU L R/O Mean Key Value Logic Overall GPT-S
Zero-shot 37.0 0.225 0.738 0.482 0.549 0.638 0.597 0.628 5.87
Fixed few-shot 42.0 0.368 0.719 0.543 0.613 0.692 0.637 0.735 6.54
Fixed few-shot + RS 52.8 0.657 0.771 0.714 0.635 0.701 0.654 0.734 6.65
Random few-shot + RS 56.5 0.727 0.791 0.759 0.626 0.731 0.670 0.772 6.74
ARAP 57.0 0.824 0.824 0.807 0.653 0.715 0.678 0.774 6.84
4.2. Analogical Prompting result on ChatGPT
4.2.1. Main Results
Table 2 displays the results of different prompting methods on ChatGPT, where ARAP represents our analogical reasoning augmented prompting. The experimental results demonstrate that: (1) ARAP significantly outperforms other prompting methods in various evaluation metrics, especially in terms of structural accuracy which improves by over 7%, proving that analogical examples provide a good logical structure for reasoning. (2) Even randomly selected few-shot reasoning examples are better than fixed ones. This indicates that fixed reasoning examples for all inputs are often suboptimal, demonstrating the necessity of establishing a reasoning library to provide diverse reasoning samples. (3) Reasoning steps are essential in prompting. When comparing results between “Fixed few-shot” and “Fixed few-shot + RS”, we found that reasoning steps have a significant active impact on reasoning accuracy, which is similar to the trend in other previous research works (Wei et al., 2022).
4.2.2. Case Study
Table 6 shows some cases we obtain from ChatGPT with different versions of prompts. Although we provide the same tag list for a specific demand across different prompt versions, there are significant variations in the selection of key-value pairs. For the first demand “Company white-collar workers who enjoy drinking Starbucks”, all baseline versions ignore the tag of “Career”, indicating that the model cannot smoothly map the description “white-collar workers” to “Career”, although this is common sense. Meanwhile, in both fixed few-shot versions, the model treats “company white-collar workers” as the value of the “Resident City”, which is due to the answers of fixed examples in prompts containing similar expressions such as (Resident city#Belongs To#Hangzhou), resulting in the model being inappropriately imitated. While in ARAP, we retrieve similar demands like “Young people who enjoy swimming or female white-collar workers who enjoy reading in Shanghai” and its reasoning steps from the reasoning library, thus helping the LLM analogically transform the “white-collar workers” into conditional expression “(Career#Belongs To#White-collar)” as they share the similar description “white-collar workers”. In another case “Young men who enjoy collecting toy models”, there are many grammar errors and inappropriate expressions in baseline answers, like non-existent operators “Contain” and inappropriate value “Young” for key “Age”. More details are described in the Appendix B.
4.3. Knowledge distillation results on smaller LLMs
Table 3. NL2SELL results on distilled model Baichuan2-13B-Chat and ChatGLM-6B-32K. The best tuning results of each fine-tuned LLMs are in bold.
Overall Acc. Structure Acc. Human Eval. GPT4 Eval.
S-BLEU L R/O Mean Key Value Logic Overall GPT-S
ChatGPT ARAP 57.0 0.653 0.824 0.807 0.815 0.715 0.678 0.774 6.84
Baichuan2-13B-Chat ARAFT 62.2 0.822 0.806 0.814 0.779 0.737 0.813 0.735 7.10
 - MT 60.8 0.809 0.794 0.802 0.754 0.731 0.800 0.718 6.98
 - RS 50.1 0.741 0.762 0.751 0.713 0.661 0.754 0.679 6.63
Normal 54.9 0.812 0.786 0.806 0.754 0.712 0.809 0.715 6.77
ChatGLM2-6B-32K ARAFT 56.0 0.829 0.773 0.801 0.740 0.715 0.790 0.710 7.13
 - MT 55.3 0.784 0.757 0.770 0.726 0.689 0.768 0.686 6.65
 - RS 55.0 0.775 0.742 0.758 0.688 0.669 0.741 0.668 6.33
Normal 42.0 0.368 0.719 0.543 0.641 0.603 0.698 0.612 5.72
4.3.1. Main Results
Table 3 shows the testing results on fine-tuned LLMs with the knowledge distilled from ChatGPT, where ARAFT represents fine-tuning (FT) in the reasoning-augmented multi-task paradigm mentioned in section 3.2.2. All testing prompts are the same as ARAP prompts in ChatGPT. As is shown: (1) Both two models show comparable capability to ChatGPT, especially on Baichuan2-13B-Chat, where the score of S-BLEU is improved by over 5% compared to ChatGPT after fine-tuning. This demonstrates the superiority of our distillation and training methods. (2) Compared with the results of the single-task training strategy (-MT), the multi-task training strategy brings improvement to the reasoning performance of the LLMs. Although during testing, we limit the model to only output the final answers without reasoning steps, adding training tasks that predict reasoning steps can enhance the models’ reasoning abilities and robustness. (3) Compared with the results without reasoning steps in the input corpus (-RS), we can conclude that providing explicit reasoning steps in training input is beneficial. If there are no reasoning steps in the input corpus but a multi-task training strategy is used, it will increase the difficulty of the training task and lead to poor performance, which is particularly evident in Baichuan2-13B-Chat. (4) The result of fine-tuning using only the normal demand and answer pairs is not satisfactory, which indicates that it is challenging for LLMs to learn the data patterns solely based on the Q&A pairs in situations where tasks are difficult. Therefore, it is necessary to design appropriate demonstrations and training tasks to fine-tune the LLMs.
Table 4. Ablation study of answer to demand distillation approach on Baichuan2-13B-Chat.
Use a2d ? Overall Acc. Structure Acc. Human Eval. GPT4 Eval.
S-BLEU L R/O Mean Key Value Logic Overall GPT-S
ARAFT 62.2 0.822 0.806 0.814 0.779 0.737 0.813 0.735 7.10
58.9 0.811 0.784 0.798 0.752 0.713 0.790 0.712 7.01
ARAFT - RS 50.1 0.741 0.762 0.751 0.713 0.661 0.754 0.679 6.63
46.1 0.667 0.728 0.697 0.674 0.625 0.700 0.637 6.37
4.3.2. Ablation Study
As mentioned in section 3.2.1, there are two sources of distilled knowledge we use when fine-tuning, i.e., knowledge distilled from demand to answer and answer to demand approach. The knowledge distilled from demand to answer is the most commonly used knowledge distillation approach, and what we mainly want to explore is the effectiveness of the knowledge distilled from answer to demand. To verify this, we conduct ablation experiments on the best-performing models “Baichuan2-13B-Chat” with two training settings (i.e., ARAFT and ARAFT - RS). Table 4 shows the effect of the knowledge distilled from answer to demand (a2d), it can be found that when removing this part of data during the fine-tuning, the performance on both ARAFT and ARAFT - RS significantly decreases, although this part of data only accounts for about one-tenth of the training data. This indicates that adding samples with a more uniform distribution of logic and tags to the training corpus can improve the robustness of the LLMs, and our distillation method from answers to demand achieves this.
5. Application
Model deployment. The LLM tuned in section 3.2 has been deployed online for application using an A10 with a memory capacity of 24G. To achieve optimal online inference performance, the deployment involves a total of 8 cards, including both the pre-production and online environments. Meanwhile, the retriever service based on BGE has also been deployed on an A10 GPU.
Application Case. Figure 3 shows an application case of ARALLM in online user targeting. When the marketers input their demand and click the “Search” button, our system will invoke the retriever service to retrieve the demands from the reasoning library most related to the current demand for generating the prompt based on analogical prompting and request the deployed ARAFT model to generate the answer expressed in SELL. The SELL expression will be parsed by our parser and results will be visualized to the marketers on the panel. The marketers can verify and edit the card on the panel (Figure 3) after clicking the “Edit” button, such as modifying the operator “AND” to “OR”. Finally, they can click the “Export” button to get the target users. Our system has been already running for months. We have set two metrics to evaluate the new approach online, including operation time, and number of likes from marketers. By collecting the marketers’ operation log and questionnaire surveys, we find the new system’s operation time is 4-10 times shorter than the former on average, and the number of likes is 1.3 times larger than the former, revealing the advantage of this new way of user targeting.
Figure 3. The application of ARALLM in User Targeting.
6. Related Work
6.1. User targeting
Methods of user targeting mainly can be categorized into two lines: rule-based and model-based methods. The rule-based methods (Mangalampalli et al., 2011; Shen et al., 2015) match potential users with specific demographic tags (age, gender, geography) or interests that are targeted by marketers which need marketers to do user profile association or mining. The model-based methods expand a given seed set by calculating the similarity of all pairs between seed users and candidate users (Ma et al., 2016; Liu et al., 2016) or training customized prediction models for each service or campaign (Zhuang et al., 2020; Zhu et al., 2021). These methods show good performance in user targeting but need seed sets. Both methods have primarily focused on intricate architectures for effectively and efficiently targeting users, while commonly disregarding the natural and significant gap between marketers’ demands and the capabilities of current models, which can be time-consuming and unfriendly.
6.2. NL2SQL
Our NL2SELL task is somewhat similar to the task of translating natural language into SQL (NL2SQL).
NL2SQL is an important and challenging task in helping non-expert users to better manipulate databases. Graph neural networks (GNNs) are popular in NL2SQL research (Wang et al., 2020; Cao et al., 2021; Hui et al., 2021), which focus on modeling the relationship between the question and the schema in database system. Recently, large pre-trained language models (PLMs), such as T5 (Raffel et al., 2019) and GPT-3 (Brown et al., 2020), have shown strong translating ability in NL2SQL tasks. Due to the extensive knowledge injection during the pretrain phase, PLM-based methods achieve better results compared to GNN-based methods by fine-tuning them with a small amount of data. Some works also explore solving NL2SQL tasks with in-context learning (Liu et al., 2023; Gu et al., 2023), it has been pointed out that LLMs show strong few-shot or zero-shot abilities even without any training data, which start a new direction for future research on NL2SQL field. The LLM-based methods in NL2SQL bring inspiration to our works.
7. Conclusion and future work
In this paper, we provide a novel user targeting approach by leveraging LLMs to gain a structured understanding of marketers’ demands. We first define a new language SELL to express the needs of marketers more clearly. Subsequently, we propose an analogical reasoning augmented framework, ARALLM, which consists of analogical reasoning based prompting and reasoning-augmented multi-task model distillation. The experimental results on ChatGPT show that our analogical reasoning based prompting significantly outperforms other baseline prompting methods in the NL2SELL task. In addition, we distill a large-scale dataset using ChatGPT and train the student LLMs using a multi-task training approach, which is successfully used for online deployment.
Since we primarily focus on the NL2SELL task in practical application scenarios, we have not yet applied the ARALLM framework to other reasoning tasks, such as coding or other structured language translation(e.g., HTML, JSON), which will be explored as future work. At the same time, there are still labor costs in the construction of reasoning libraries and SELL datasets, we will explore more automated construction methods in the future.
Appendix A Logic template
Table 5 shows the logic template we use for answer generation in section 3.2.1. The c in the template represents the basic conditional expression in SELL.
Table 5. Logic template for answer generation.
Template
c
c AND c
c AND c AND c
c AND c AND c AND c
c OR c
c OR c OR c
c OR c OR c OR c
(c AND c) OR c
(c OR c) AND c
(c AND c) OR (c AND c)
(c OR c) AND (c OR c)
c AND (c OR c OR c)
c OR (c AND c AND c)
(c AND c) OR (c OR c)
(c AND c) AND (c OR c)
c AND ((c AND c) OR c)
c AND ((c OR c) AND c)
c OR ((c OR c) AND c)
c OR ((c AND c) OR c)
Table 6. Case study on ChatGPT with different versions of prompt. Bad responses are marked with a wavy underline.
Demand:Company white-collar workers who enjoy drinking Starbucks
Zero-shot
(Preference#Belongs To#Starbucks)
Fixed few-shot
(Preference#Belongs To#Starbucks) AND (Resident City#Belongs To#Company white-collar)
Fixed few-shot + RS
(Preference#Belongs To#Enjoy drinking Starbucks) AND (Resident City#Belongs To#Company white-collar)
Random few-shot + RS
(Preference#Belongs To#Starbucks)
ARAP (Preference#Belongs To#Starbucks) AND (Career#Belongs To#White-collar)
Demand:Young men who enjoy collecting toy models
Zero-shot
(Preference#Contain#Toy models) AND (Gender#Equal To#Male) AND (Age#Belongs To#[18,30])
Fixed few-shot
(Preference#Belongs To#Toy models) AND (Gender#Belongs To#Male) AND (Age#Less Than#30)
Fixed few-shot + RS
(Preference#Belongs To#Toy models) AND (Gender#Belongs To#Male)
Random few-shot + RS
(Preference#Belongs To#Toy models) AND (Gender#Belongs To#Male) AND (Age#Belongs To#Young)
ARAP (Preference#Belongs To#Toy models) AND (Gender#Belongs To#Male) AND (Age#Between#18,35)
Appendix B Case study on ChatGPT
Table 6 shows two cases we obtain from ChatGPT with different versions of prompts. For “Fixed few-shot” and “Fixed few-shot + RS”, we sample two examples from reasoning library $\mathcal{R}$. These two fixed demands are: “Women who like Estee Lauder and live in Hangzhou” and “Users who have a certain understanding or interest in financial products or insurance”. Their corresponding SELL answers are “(Preference#Belongs To#Estee Lauder) OR (Luxury Fashion Preference User#Belongs To#True) AND (Resident City#Belongs To#Hangzhou City) AND (Gender#Belongs To#Female)” and “(Preference#Belongs To#Financial Products) OR (Preference#Belongs To #Insurance)”. For ARAP, we retrieve two similar demands in $\mathcal{R}$ for each testing demand based on their textual embedding similarity. For the first case, we obtain “Young people who enjoy swimming or female white-collar workers who enjoy reading in Shanghai” and “Young white-collar workers in Jianghan District, Wuhan City”. For the second case, we obtain “Young people who like high-end cosmetics” and “Young Readers Exploring the World of Science Fiction”.
Appendix C Demand generation prompt
The prompt framework for demand generation given SELL expressions is shown in Figure 4.
Figure 4. Demand generation prompt.
Appendix D GPT4 Evaluation Prompt
The prompt framework for guiding GPT4 to evaluate the quality of answer SELL is shown in Figure 5.
Figure 5. GPT4 Evaluation prompt.